OpenStudy (anonymous):

Please help, i will fan & medal!! no one is helping me The legs of a right triangle are 3 units and 6 units. What is the length of the hypotenuse? Round your answer to the nearest hundredth. i know this part: a^2 + b^2 = c^2 6^2 + 3^2 36 + 9 = ?

OpenStudy (texaschic101):

so what is 36 + 9 ?

OpenStudy (anonymous):

i know that it's 45 but it's not an answer choice

OpenStudy (texaschic101):

a^2 + b^2 = c^2 6^2 + 3^2 = c^2 36 + 9 = c^2 45 = c^2 sq rt 45 = c 6.708 = c

OpenStudy (anonymous):

ohhhh i forgot square root

OpenStudy (texaschic101):

to get rid of ^2, you use the square root

OpenStudy (anonymous):

thank you so much!

OpenStudy (texaschic101):

sure thing :)